OpenStudy (anonymous):

Hi i am just going over the investment word problems in purple math...does anyone have th answers for the additional examples?

OpenStudy (anonymous):

post the problem on which you want clarification.

OpenStudy (anonymous):

An investment of $3,000 is made at an annual simple interest rate of 5%. How much additional money must be invested at an annual simple interest rate of 9% so that the total annual interest earned is 7.5% of the total investment?

OpenStudy (anonymous):

"annual simple interest rate" means we can use rate * time* principal=amount formula (x*.09*1)+(3,000 * .05*1)=(3000+x)*(.075)

OpenStudy (anonymous):

solve for x

OpenStudy (anonymous):

i got 500....

OpenStudy (anonymous):

close , it is 5000

OpenStudy (anonymous):

(x*.09*1)+(3,000 * .05*1)=(3000+x)*(.075) .09x+150=225+0.075x

OpenStudy (anonymous):

oh yeah i see the mistake....thank you!

OpenStudy (anonymous):

my mistake

OpenStudy (anonymous):

thank you!